# When Should Selfish Miners Double-Spend?

### v1

@@ 3.1 $L$-Stubborn Mining
The selfish mining strategies are extensively analyzed in the literature with notable works such as [4, 7, 6]. In this work, we consider a sub-group of strategies from all possible states and action space, which we call $L$-stubborn [^1] In this paper, $L$-stubborn mining is different than the lead-stubbornness (L-stubbornness) defined in [7]. There, “L” stands for the word “lead.” Here, $L$ refers to an integer which is the length of the adversarial chain with respect to a common ancestor block between the honest and the adversarial chains. mining. To describe the strategy, let us assume that all participants start with the same chain, which we call the offset chain. Both the adversary and the honest miners mine on top of the offset chain. If the honest miners are the first to mine, then the adversary accepts the new block and redefines the offset chain with the new block on its tip. For the sake of simplicity, we denote the height of the offset chain as zero. If the adversary is the first to mine a block on top of the offset chain, i.e., at height $1$, it keeps the block private and we denote the highest block height of the adversarial chain as $A$ and call the adversarial chain as $A$-chain. At any time later, if the honest miners mine a block on that height, adversary publishes its block on that height immediately if $A<L$, which we call matching. In accordance with the network model defined earlier, $\gamma$ proportion of the honest miners accept and mine on top of the adversarial block instead of the honest block. We denote the highest honest block height as $H$ and the chain starting from the offset and ending in the honest block at height $H$ as $H$-chain. The private stubborn mining strategy of the adversary continues on for all heights until $A=L$, i.e., whenever $L>A\geq H$ adversary immediately matches the block at the height of the newly mined honest block. If honest miners manage to mine a block on a height before the adversary before $A=L$, i.e., $L\geq H=A+1$, the adversary accepts the $H$-chain and redefines the offset chain. On the other hand if $A\geq L$, then the adversary keeps mining privately until $A=H+1$, at which point, it releases the $A$-chain and redefines the offset chain. An attack cycle is a process that starts when an offset chain is defined, and ends when a new offset chain is defined.
@@
Note that, for $L=1$ the strategy reduces to honest mining, whereas for $L=2$ the strategy reduces to the celebrated selfish mining attack of [4]. On the other hand, for $L=\infty$ the strategy becomes the equal fork stubborn mining strategy described in [7] and analyzed in closed-form with Catalan numbers in [28]. In this paper, we are interested in giving a method for calculating the revenues for any given $L$ as well as deriving a method to find the best possible $L$ given $\alpha$ and $\gamma$. The resulting strategy does not allow any trail stubbornness discussed in [7], i.e., the adversary accepts the honest chain whenever it falls behind. Hence, we can only claim the optimality of the strategy under action spaces with independent attack cycles that allow no-trail stubbornness. We argue that the parameter $L$ represents the level of stubbornness in selfish mining and our analysis provides the optimal level of stubbornness as well as the maximum stubbornness such that the attack is still more profitable than honest mining. Thus, we view $L$ (and $S$, which will be defined later) as a parameter of stubbornness.
@@
In [6], the authors use an undiscounted average reward MDP to provide $\epsilon$-optimal policies. $L$-stubborn mining allows a sub-group of strategies from all possible states and action space and we give an easy-to-evaluate formula to find optimal $L$. The authors of [6] further argue about combining double-spending attack with selfish mining attack, however, they do not provide any concrete method or analysis. Later, in [9], the same authors provide an MDP that accounts for rewards of double-spending attack, however, the MDP ignores rewards from adversarial blocks that do not result in double-spending. Thus, the optimal policies resulting from the MDP therein are not optimal in the sense that they provide maximum mining rewards with free double-spending. Further, the results and strategies provided in [9, 6, 8] are complicated with no easy-to-follow structure. $L$-stubborn mining that we consider in this paper is a policy which results in higher mining rewards than honest mining (as well as optimal ones) and has strong implications on double-spending strategies with a clear and easy-to-follow analysis under $k$-confirmation rule, which we explain next.
@@ 3.2 $k$-Confirmation Rule
Before the offset event happens, if $A>k$, this implies that at some point $H=k$ and adversary will release $A$-chain later on. Depending on whether $H$-chain contains $tx^{\prime}$ or $tx$, either Move-Funds or Double-Spending happens. On the other hand, before the offset event happens, if $A<H\leq k$, depending on whether $H$-chain contains $tx^{\prime}$ or $tx$, either Move-Funds or Service happens. If an attack cycle ends with Move-Funds event, later, adversary can try double-spending again using the same funds. This also decreases effective throughput in finite block size systems. One could also rescale the probabilities of Double-Spending and Service by excluding Move-Funds event from the perspective of the actual funds associated with $tx$ and $tx^{\prime}$ since eventually the fund is going to result in Service or Double-Spending. We omit the rescaling in our calculations and numerical results, as rescaling is straightforward to do and keeping Move-Funds event gives an understanding of throughput decrease to the reader.
We also note that the time $tx$ is released in the network is another important issue while considering double-spending attacks. In general, it can be shown that a double-spending attack can be successful with probability $1$ if adversary is allowed to pick the time $tx$ is released, since it can build an arbitrarily high pre-mining gain [9]. The pre-mining gain refers to the difference between the private chain of the adversary and the public honest chain, in general [9]. The approach we take here assumes that the $tx$ is released at the start of each attack cycle, hence the pre-mining gain is 00. Thus, the results we provide are conservative and the adversary could get even more profits from double-spending with more clever approaches. Indeed, as we will see later in the numerical results section, the service probability is always above $\beta$, since the first block mined by the honest nodes contains $tx$ and with probability $\beta$ it is mined before any adversarial blocks and immediately becomes part of the offset chain.
@@ 3.3 $S$-Stealth Mining
We address some of these issues by modifying $L$-stubborn mining in a way that only allows matching if both double-spending and the risk of losing mining revenue are imminent, i.e., $H=A=k=L-1$. We call this attack $S$-stealth mining, which we investigate in Section 5. We note that $S$-stealth mining does not allow any trailing behind either, hence, there is still a difference between a private attack and $S$-stealth mining.
4 Revenue & Double-Spending of $L$-Stubborn Mining
@@
Instead of building a complex Markov chain and going through the tedious analysis of steady-state distribution, we can analyze the distinct possible paths an attack cycle can take. An attack cycle starts when an offset chain is defined, and ends when a new offset chain is defined. If during an attack cycle, the length of the $A$-chain reaches $L$, i.e., $A=L$, then we call the attack successful. On the other hand, if the length of the $H$-chain surpasses the length of the $A$-chain before $A$ reaches $L$, we call the attack cycle unsuccessful.
@@
In Figure 1, we display the sample paths of $4$-stubborn mining, i.e., $L=4$. The $y$-axis and $x$-axis represent $A$ and $H$, i.e., the length of $A$-chain and $H$-chain, respectively. An attack starts from point $(0,0)$ and arrivals of adversarial (honest resp.) blocks are represented with light-green (orange resp.) arrows. A red arrow shows that the arrival of honest block has ended the cycle. A blue (red resp.) cross represents the endpoint of a successful (unsuccessful resp.) attack cycle. A dark-green arrow represents the arrival of an adversarial block that guarantees a successful attack cycle. The dark-green crosses represent the first time cycles are guaranteed to be successful and not ended yet. The cycles that pass through the dark-green crosses eventually end with an honest arrival in blue cross points $(x,x-1)$ where $x>4$, i.e., when the difference of chain lengths reduces to $1$.
@@
Note that there is a third dimension of Figure 1, which tracks the number of adversarial blocks in the prefix of $H$-chain since every arrival of honest block has two choices of chains to extend. Further, the dynamics of this dimension is Markovian since each honest arrival either increases the previous value to $H-1$ (w.p. $\gamma$) or stays the same (w.p. $1-\gamma$) and does not depend on anything else. This Markovian structure implies that we can separate the third dimension and treat it independently which simplifies our analysis, hence, we do not plot it.
@@
Under the $L$-stubborn mining strategy, the probability that $A$-chain reaches length $L$ during an attack cycle while the largest honest block is at height $m<L$ is
@@
Proof:  We know that the winner of the cycle is determined once the last adversarial block arrives, hence we distinguish it from the previous $L-1$ blocks to prevent double-counting. The number of distinct paths with $H=m<L$ blocks and $A=L$ blocks, where the last block is adversarial, is equal to $P[L-1,m]$; see Figure 1 for an example where $L=4$. $\blacksquare$
@@
The probability that an attack cycle is unsuccessful and ends with the largest honest block at height $n+1$ which has $i\leq n$ adversarial blocks in its prefix (after the offset chain) is
@@
Further, each time an honest block arrives, it is either mined by the honest miners who are under the influence of adversarial view w.p. $\gamma$, or it is mined on top of the last honest block. As a result, no adversarial block is contained in the prefix of the honest block on height $n+1$ w.p. $(1-\gamma)^{n}$ or $i>0$ adversarial blocks are contained w.p. $(1-\gamma)^{n-i}\gamma$. Note that, $\mathbbm{1}_{i}=1$ if $i\neq 0$ and zero otherwise. $\blacksquare$
Next, we put Lemma 1 and Lemma 2 together to obtain the revenue of selfish mining.
@@
(Revenue) The revenue of $L$-stubborn mining is given as
@@
Proof:  In Lemma 1 and Lemma 2, we provide the probability for each possible outcome of an attack cycle. If an attack cycle is successful, we have an $A$-chain at length $L$ and the largest honest block is at height $m<L$. Clearly, if $L-m>1$, the adversary will continue the cycle until the difference goes down to $1$. Using Wald’s identity provided in Section 2.3, the expected revenue in this case is $L+\frac{(L-m-1)\alpha}{1-2\alpha}$. This explains $L_{s}$ in (10). On the other hand, if the attack cycle is unsuccessful, $i$ adversarial blocks still make it into the offset chain w.p. $P_{u}(n,i)$ and the total number of blocks that make it into the offset chain is $n+1$. As a result, these two cases are considered in the numerator and denominator of (10) as $L_{u,a}$ and $L_{u}$, respectively. $\blacksquare$
Note that, a successful attack cycle does not immediately imply a successful double-spending. It only implies that the $A$-chain will make it into the offset chain eventually. Similarly, an unsuccessful attack cycle does not immediately imply that $tx$ will be confirmed.
Next, we provide the probability of each event regarding $tx$ and $tx^{\prime}$ for $(k+1)$-stubborn mining with $k$-confirmation rule.
Theorem 2
[…]
$P(\textbf{Service})$ $=\sum_{m=0}^{k}P_{u}(m)(1-\gamma)^{m},$ (15)
[…]
4.1 Double-Spending Danger is Imminent Well Before $\alpha=0.5$
We define
$L^{*}$ $=\arg\max_{L}\rho_{L},$ (17)
$\bar{L}$ $=\sup\{L:\rho_{L}\geq\alpha\},$ (18)
where $\bar{L}\geq L^{*}$ due to the quasiconcavity proven in Theorem 3. Essentially, any $k$-confirmation rule with $k<\bar{L}$ is compromised by $(k+1)$-stubborn mining since adversary gains more revenues than honest mining, i.e., $\rho_{k+1}\geq\alpha$, and at the same time it can double-spend at no-cost. In other words, for a given $\alpha$ and $\gamma$, $k$-confirmation rule should be at least $k\geq\bar{L}$ to make sure that double-spending attack has revenue loss. Note that, if $\rho_{\infty}\geq\alpha$, no $k$-confirmation is safe as double-spending has zero-cost. From this perspective, the safety of a blockchain system fails well before adversarial proportion reaches $\alpha$. We investigate this phenomenon further in the numerical results section by finding $L^{*}$ and $\bar{L}$ for each $\alpha$ and $\gamma$.

### v2

@@ 3.2 $L$-Stubborn Mining
The selfish mining strategies are extensively analyzed in the literature with notable works such as [4, 6, 5]. In this work, we consider a sub-group of strategies from all possible states and action space, which we call $L$-stubborn [^1] In this paper, $L$-stubborn mining is different than the lead-stubbornness (L-stubbornness) defined in [6]. There, “L” stands for the word “lead.” Here, $L$ refers to an integer which is the length of the adversarial chain with respect to a common ancestor block between the honest and the adversarial chains. mining. To describe the strategy, let us assume that all participants start with the same chain, which we call the offset chain. For the sake of simplicity, we denote the height of the offset chain as zero. Both the adversary and the honest miners mine on top of the offset chain.
  1. 1.
    If the honest miners are the first to mine, then the adversary accepts the new block and redefines the offset chain with the new block on its tip.
  2. 2.
    If the adversary is the first to mine a block on top of the offset chain, i.e., at height $1$, it keeps the block private and continues to mine privately on top of its block. We denote the highest block height of the adversarial chain as $A$ and call the adversarial chain as $A$-chain. At any time later, if the honest miners mine a block on at height $h$, the adversary publishes its block at the same height immediately if $h\leq A<L$, which we call matching. In accordance with the network model defined earlier, $\gamma$ proportion of the honest miners accept and mine on top of the adversarial block instead of the honest block. We denote the highest honest block height as $H$ and the chain starting from the offset and ending in the honest block at height $H$ as $H$-chain. The private stubborn mining strategy of the adversary continues on for all heights until $A=L$, i.e., whenever $L>A\geq H$ adversary immediately matches the block at the height of the newly mined honest block.
    1. (a)
      If honest miners manage to mine a block on a height before the adversary before $A=L$, i.e., $L\geq H=A+1$, the adversary accepts the $H$-chain and redefines the offset chain. Such an action of accepting $H$-chain and abandoning $A$-chain is called adopt.
    2. (b)
      If honest miners do not manage to mine a block on a height before the adversary and $A$ chain reaches the length $A=L$, then the adversary keeps mining privately until $A=H+1$, at which point, it releases the $A$-chain and redefines the offset chain. Such an action of releasing an $A$-chain with $A>H$ is called override.
An attack cycle is a process that starts when an offset chain is defined, and ends when a new offset chain is defined. Note that, offset chain is redefined whenever an adopt or override action (event) happens, which we call an offset event in general.
@@
Note that, for $L=1$ the strategy reduces to honest mining strategy, whereas for $L=2$ the strategy reduces to the celebrated selfish mining attack of [4]. On the other hand, for $L=\infty$ the strategy becomes the equal fork stubborn mining strategy described in [6] and analyzed in closed-form with Catalan numbers in [37]. In this paper, we are interested in giving a method for calculating the revenue ratios for any given $L$ as well as deriving a method to find the best possible $L$ given $\alpha$ and $\gamma$. The resulting strategy does not allow any trail stubbornness discussed in [6], i.e., the adversary accepts the honest chain whenever it falls behind. Hence, we can only claim the optimality of the strategy under action spaces with independent attack cycles that allow no-trail stubbornness. We argue that the parameter $L$ represents the level of stubbornness in selfish mining and our analysis provides the optimal level of stubbornness as well as the maximum stubbornness such that the attack is still more profitable than honest mining.
@@
In [5], the authors use an undiscounted average reward MDP to provide $\epsilon$-optimal policies. The authors of [5] further argue about combining double-spending attack with selfish mining attack, however, they do not provide any concrete method or analysis. Later, in [8], the same authors provide an MDP that accounts for rewards of double-spending attack, however, the MDP ignores rewards from adversarial blocks that do not result in double-spending. Thus, the optimal policies resulting from the MDP therein are not optimal in the sense that they provide maximum mining rewards with free double-spending. Further, the results and strategies provided in [8, 5, 7] are complicated with no easy-to-follow structure. $L$-stubborn mining that we consider in this paper is a policy which results in higher mining rewards than honest mining (as well as optimal ones) and has strong implications on double-spending strategies with a clear and easy-to-follow analysis under $k$-confirmation rule, which we explain next.
@@ 3.3 $k$-Confirmation Rule
Before the offset event happens, if $A>k$, this implies that at some point $H=k$ and adversary will release $A$-chain later on. Depending on whether $H$-chain contains $tx^{\prime}$ or $tx$, either Move-Funds or Double-Spending happens. On the other hand, before the offset event happens, if $A<H\leq k$, depending on whether $H$-chain contains $tx^{\prime}$ or $tx$, either Move-Funds or Service happens. If an attack cycle ends with Move-Funds event, later, adversary can try double-spending again using the same funds. One could also rescale the probabilities of Double-Spending and Service by excluding Move-Funds event from the perspective of the actual funds associated with $tx$ and $tx^{\prime}$ since eventually the fund is going to result in Service or Double-Spending. We omit the rescaling in our calculations and numerical results, as rescaling is straightforward to do and we would like to present the exact probability of Double-Spending event in each attack cycle.
We also note that the time $tx$ is released in the network is another important issue while considering double-spending attacks. In general, it can be shown that a double-spending attack can be successful with probability $1$ if the adversary is allowed to pick the time $tx$ is released, since it can build an arbitrarily high pre-mining gain [8]. The pre-mining gain refers to the difference between the private chain of the adversary and the public honest chain, in general [8]. The approach we take here is more conservative and assumes that the $tx$ is released at the start of each attack cycle, hence the pre-mining gain is 0.
@@ 3.4 $S$-Stealth Mining
We address some of these issues by modifying $L$-stubborn mining in a way that only allows matching if both double-spending and the risk of losing mining revenue are imminent, i.e., H=A=k=L1H=A=k=L-1. We call this attack $S$-stealth mining, which we investigate in Section 5. Essentially, $S$-stealth mining strategy is the same strategy as $L$-stubborn mining strategy except that the matching action only happens when H=A=S1H=A=S-1. We note that $S$-stealth mining does not allow any trailing behind either, hence, there is still a difference between a private attack and $S$-stealth mining. Note that, similar to $L$ of $L$-stubborn mining strategy, we can view $S$ of $S$-stealth mining strategy as a parameter of stubbornness.
3.5 Definitions and Notations
In this section, we give definitions and notations that are used frequently throughout the paper. For the convenience of the reader, some definitions provided in the other parts of the paper are listed here as well.
Definition 2
(Model parameters) The fraction of the total hashrate the attacker pool possesses is denoted as $\bm{\alpha}$. The remaining miners are considered to be honest and possess $\bm{\beta}=1-\alpha$ fraction of the hashrate. When a public fork exists, $\bm{\gamma}$ fraction of the honest miners prefer the adversarial branch.
Definition 3
($\bm{k}$-Confirmation rule) A block is said to be confirmed according to the $k$-confirmation rule if it is part of the longest public chain and $k$-deep, i.e., there are at least k1k-1 public blocks mined on top of it in the longest chain.
Definition 4
(Mining strategies) The following strategies are frequently mentioned in the paper:
  1. 1.
    $\bm{L}$-stubborn mining strategy refers to the mining strategy for which the procedure is explicitly explained in Section 3.2.
  2. 2.
    Honest mining strategy refers to the mining strategy when the pool strictly follows the honest protocol and publishes a block as soon as it is mined. The strategy can be explicitly obtained by plugging $\bm{L=1}$ into the procedure explained in Section 3.2.
  3. 3.
    Selfish mining strategy refers to the mining strategy proposed by Eyal and Sirer [4]. The strategy can be explicitly obtained by plugging $\bm{L=2}$ into the procedure explained in Section 3.2. We also give an explicit summary of the attack in Appendix A as well as how $L$-stubborn mining generalizes the attack.
  4. 4.
    Equal fork stubborn mining strategy refers to the mining strategy proposed by [6]. The strategy can be explicitly obtained by plugging $\bm{L=\infty}$ into the procedure explained in Section 3.2.
  5. 5.
    $\bm{S}$-stealth mining strategy refers to the mining strategy for which the procedure is explicitly explained in Section 3.4.
Definition 5
(Revenue ratio) Revenue ratio of the pool refers to the fraction of the blocks created by the pool in the longest chain in the long run. In other words,
$\bm{\rho=\frac{\#blocks~created~by~pool}{\#total~blocks~in~longest~chain}}$ (7)
Here,
  1. 1.
    $\bm{\rho_{L}}$ is the revenue ratio of $L$-stubborn mining strategy, for which an explicit equation is provided in Theorem 1, Eq. (18). Here,
    1. (a)
      $\bm{\rho_{1}}$ is the revenue ratio when the pool employs the honest mining strategy, where
      $\rho_{1}=\alpha.$ (8)
    2. (b)
      $\bm{\rho_{2}}$ is the revenue ratio of selfish mining strategy, where
      $\rho_{2}=\frac{\alpha\beta^{2}(4\alpha+\gamma(1-2\alpha))-\alpha^{3}}{1-\alpha(1+(2-\alpha)\alpha)}.$ (9)
    3. (c)
      $\bm{\rho_{\infty}}$ is the revenue ratio of equal fork stubborn strategy, where
      $\rho_{\infty}$ $=\frac{\alpha}{\beta}-\frac{(1-2\alpha)(1-\gamma)}{\beta\gamma}(1-\beta C((1-\gamma)\alpha\beta)).$ (10)
  2. 2.
    $\bm{\sigma_{S}}$ is the revenue ratio of $S$-stealth mining strategy, for which an explicit equation is provided in Theorem 5, Eq. (31).
Remark 1
In the long run, for every coin issued by the protocol, the pool gets $\rho$ fraction of it. Note that, for a given $(\alpha,\gamma)$ pair, if a specific mining strategy by the attacker pool results in a revenue ratio higher than the revenue ratio of honest mining strategy, i.e., $\rho>\alpha$, the system is not incentive compatible for the $(\alpha,\gamma)$ pair.
Remark 2
Some mining strategies potentially result in double-spending profits in addition to the revenues from the coinbase rewards of the blocks. These additional profits are not considered in revenue ratio as it is not the convention in selfish mining literature. Instead, we define a combined revenue ratio in Section 4.3 that accounts for the additional profit coming from the double-spending similar to the subversion gain defined in [10].
Remark 3
Normalized revenue ratio is defined as $\frac{\rho}{\alpha}$, which is a measure of how much the revenues are scaled when a miner employs a non-honest mining strategy compared to the honest mining strategy.
Definition 6
$L^{*}$ refers to the value of $L$ that maximizes the revenue ratio of $L$-stubborn mining, i.e.,
$L^{*}$ $=\arg\max_{L}\rho_{L}.$ (11)
Further, $\bar{L}$ refers to the largest value of $L$ such that the revenue ratio of $L$-stubborn mining is higher than the honest mining strategy,
$\bar{L}$ $=\sup\{L:\rho_{L}\geq\alpha\}.$ (12)
Similarly, for $S$-stealth mining, we define
$S^{*}$ $=\arg\max_{S}\sigma_{S},$ (13)
$\bar{S}$ $=\sup\{S:\sigma_{S}\geq\alpha\}.$ (14)
We note that the following are defined in Section 3.2: $L$, AA, $H$, AA-chain, $H$-chain, offset chain, offset event, matching, override, adopt actions. As it is not possible to define them without completely going through the algorithm of $L$-stubborn mining, we refer the reader to the first paragraph of Section 3.2 where they are defined.
4 Analysis of $L$-Stubborn Mining
@@
We start by analysis of the revenue ratio of the attacker pool when it employs the $L$-stubborn mining strategy. Instead of building a complex Markov chain and going through the tedious analysis of steady-state distribution as it is done in [4], we can analyze the distinct possible paths an attack cycle can take. An attack cycle starts when an offset chain is defined, and ends when a new offset chain is defined. If during an attack cycle, the length of the $A$-chain reaches $L$, i.e., $A=L$, then we call the attack successful. On the other hand, if the length of the $H$-chain surpasses the length of the $A$-chain before $A$ reaches $L$, we call the attack cycle unsuccessful.
@@
In Figure 1, we display the sample paths of $4$-stubborn mining, i.e., $L=4$. The $y$-axis and $x$-axis represent $A$ and $H$, i.e., the length of $A$-chain and $H$-chain, respectively. An attack starts from point $(0,0)$ and arrivals of adversarial (honest resp.) blocks are represented with light-green (orange resp.) arrows. A red arrow shows that the arrival of honest block has ended the cycle. A blue (red resp.) cross represents the endpoint of a successful (unsuccessful resp.) attack cycle. A dark-green arrow represents the arrival of an adversarial block that guarantees a successful attack cycle. The dark-green crosses represent the first time cycles are guaranteed to be successful and not ended yet. The cycles that pass through the dark-green crosses eventually end with an honest arrival in blue cross points (x,x1)(x,x-1) where $x>4$, i.e., when the difference of chain lengths reduces to $1$.
@@
Note that there is a third dimension of Figure 1, which tracks the number of adversarial blocks in the prefix of $H$-chain since every arrival of honest block has two choices of chains to extend. Further, the dynamics of this dimension is Markovian since each honest arrival either increases the previous value to H1H-1 (with probability $\gamma$) or stays the same (with probability $1-\gamma$) and does not depend on anything else. This Markovian structure implies that we can separate the third dimension and treat it independently which simplifies our analysis, hence, we do not plot it.
@@
Under the $L$-stubborn mining strategy, the probability that $A$-chain reaches length $L$ during an attack cycle while the maximum height of the honest block is $H=m<L$, is
@@
Proof:  We distinguish the last adversarial block from the previous L1L-1 adversarial blocks to prevent double-counting. The number of distinct paths with $H=m<L$ blocks and $A=L$ blocks, where the last block is adversarial, is equal to P[L1,m]P[L-1,m]; see Figure 1 for an example where $L=4$.  $\blacksquare$
@@
The probability that an attack cycle is unsuccessful and ends with $H=n+1$ which has $i\leq n$ adversarial blocks in its prefix (after the offset chain) is
@@
Further, each time an honest block arrives, it is either mined by the honest miners who are under the influence of adversarial view with probability (w.p.) $\gamma$, or it is mined on top of the last honest block. As a result, no adversarial block is contained in the prefix of the honest block on height $n+1$ w.p. $(1-\gamma)^{n}$ or $i>0$ adversarial blocks are contained w.p. $(1-\gamma)^{n-i}\gamma$. Note that, $\mathbbm{1}_{i}=1$ if $i\neq 0$ and zero otherwise.  $\blacksquare$
Next, we put Lemma 1 and Lemma 2 together to obtain the revenue ratio of selfish mining.
@@
(Revenue Ratio) The revenue ratio of $L$-stubborn mining is given as
@@
Proof:  First note that, in the long run, the longest chain is essentially the offset chain that grows when an adopt or override happens. As a result, the revenue ratio of the attacker is the fraction of the blocks created by the attacker in the offset chain. As the offset event is a stopping time, we can simply focus on a single attack cycle and calculate the expected number of blocks created by the adversary that become part of the offset chain. Dividing this number to the expected growth of the offset chain in an attack cycle gives us the revenue ratio. In Lemma 1 and Lemma 2, we provide the probability for each possible outcome of an attack cycle. If an attack cycle is successful, we have an $A$-chain at length $L$ and the maximum height of the honest block is $H=m<L$ at some time within the attack cycle. Clearly, if Lm>1L-m>1, the adversary will continue the cycle until the difference goes down to $1$. Thus, given the attack is successful, using Wald’s identity provided in Section 2.3, the expected number of blocks that become part of the offset chain in this case is $L+\frac{(L-m-1)\alpha}{1-2\alpha}$ and all these blocks are adversarial. This explains $L_{s}$ in (18). On the other hand, if the attack cycle is unsuccessful, $i$ adversarial blocks still make it into the offset chain w.p. $P_{u}(n,i)$ and the total number of blocks that make it into the offset chain is $n+1$. As a result, these two cases are considered in the numerator and denominator of (18) as $L_{u,a}$ and $L_{u}$, respectively.  $\blacksquare$
4.1 Optimizing the Revenue Ratio